OpenStudy (anonymous):

easy question Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 + 10x + 25 In which direction and by how many units should f(x) be shifted to obtain g(x)? Right by 5 units Left by 5 units Right by 10 units Left by 10 units

OpenStudy (blank ):

f(x) should be shifted "Left by 5 units" to get the same results as g(x)

OpenStudy (anonymous):

howd u get the answer

OpenStudy (blank ):

You just graph them and see

OpenStudy (anonymous):

oh okay can you help with another I will post another question

OpenStudy (blank ):

Sure

OpenStudy (arbershabani97):

y = a(x – h)2 + k so the second graph is g(x)=(x+5)^2 ,so left 5 units

OpenStudy (blank ):

Thanks @arbershabani97

OpenStudy (arbershabani97):

you're welcome